OpenStudy (anonymous):

8000 deposited at 5% compounded quarterly to reach atleast 16000

OpenStudy (anonymous):

@myininaya @amistre64

OpenStudy (anonymous):

@Zarkon

OpenStudy (amistre64):

well, we start with a balance of 0, and add payments quarterly or, we could just start with a payment as the balance \[B_n=Pk^n+P\frac{k^n-1}{k-1}\] and solve for n

OpenStudy (amistre64):

isnt 8000 * 2 = 16000? just wondering

OpenStudy (amistre64):

so thats what, 3 months?

OpenStudy (amistre64):

\[16000=8000(1+\frac{k^n-1}{k-1})\] \[2=1+\frac{k^n-1}{k-1}\] \[1=\frac{k^n-1}{k-1}\] \[k-1=k^n-1\] \[k=k^n\] \[n=1\] so, 1 compounding period is a quarter of a year